OpenStudy (anonymous):

Which of the following shows the square completed in the this equation? x2 + 7x + 2 = 0 (x = 7)2 = 2 (x + 3.5)2 = 10.25 (x + 2.6)2 = 4.76

OpenStudy (blockcolder):

\[x^2+7x=-2\\ x^2+7x+\left ( \frac{7}{2}\right )^2=-2+\left ( \frac{7}{2}\right )^2\\ \left ( x+\frac{7}{2}\right )^2=-2+\frac{49}{4}\\ (x+3.5)^2=10.25\]

OpenStudy (anonymous):

thank you!

OpenStudy (blockcolder):

No prob. :)